OpenStudy (anonymous):

how would you simplify or solve the expression 5^3 · 3^2) ?

jhonyy9 (jhonyy9):

your idea ?

jhonyy9 (jhonyy9):

or this will help you ? 5*5^2 *3^2 =

OpenStudy (anonymous):

i dont understand

jhonyy9 (jhonyy9):

e you know that a^2 = a*a yes ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

or a^2 *a^3 = a^(2+3) = a^5

jhonyy9 (jhonyy9):

do you understand it ?

OpenStudy (anonymous):

so would it b 8x5

jhonyy9 (jhonyy9):

why 8 ?

OpenStudy (anonymous):

5^3 or is it 15

jhonyy9 (jhonyy9):

why 15 so 5^3 = 5*5*5 =

OpenStudy (anonymous):

25

OpenStudy (anonymous):

@jhonyy9 im confused

jhonyy9 (jhonyy9):

so than us this a^2 * b^2 = (a*b)^2

OpenStudy (anonymous):

or is 125 x 27

jhonyy9 (jhonyy9):

so i have wrote there that 5^3 = 5*5^2 and there is again 3^2 so 5*5^2 *3^2 = 5*(5*3)^2 = 5*15^2 =

jhonyy9 (jhonyy9):

can you calculi it now ?

jhonyy9 (jhonyy9):

15^2 =15*15 = 225 5*15^2 =5*225=

OpenStudy (anonymous):

1125

jhonyy9 (jhonyy9):

yes but do you understand it now how can you simplifie it and calculi easy ?

OpenStudy (anonymous):

i think so, 125 x 27 is tht right

OpenStudy (anonymous):

3375

jhonyy9 (jhonyy9):

how you have got 27 ?

jhonyy9 (jhonyy9):

because 3^2 =3*3 or not ?

jhonyy9 (jhonyy9):

@Avon do you understand it ?

jhonyy9 (jhonyy9):

there 5 is cubed just and 3 is squared

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

125 x 9=1125

jhonyy9 (jhonyy9):

yes this was easy now ok ?

jhonyy9 (jhonyy9):

ok yw bye